OpenStudy (anonymous):

Product of two summations? Say I am multiplying two polynomials: \[ \left( \sum_{k=0}^na_kx^k\right) \left( \sum_{j=0}^mb_jx^j\right)\] apparently this is equal to \[ \sum_{r=0}^{n+m}\left(\sum_{k=0}^ra_kb_{r-k}\right)x^r\]

OpenStudy (anonymous):

What I did: \[ \begin{align} \left( \sum_{k=0}^na_kx^k\right) \left( \sum_{j=0}^mb_jx^j\right) &= \sum_{k=0}^n\sum_{j=0}^ma_kb_jx^{k+j}\\ &=\sum_{k=0}^n\sum_{r-k=0}^ma_kb_{r-k}x^r \\ &=\sum_{k=0}^n\sum_{r=k}^ma_kb_{r-k}x^r \end{align}\] Obviously I'm having trouble transitioning to the required double sum with the required indices. I tried drawing a Cartesian diagram in both cases using \(m=5\) and \(n=2\) as an example to see what points are covered by the sum, but they don't seem to be the same? Any help would be appreciated.

OpenStudy (anonymous):

Oh yeah I let \(r=j+k\)

OpenStudy (anonymous):

Do you still need help?

OpenStudy (ikram002p):

why u let r=k-j ?

OpenStudy (ikram002p):

that change second summation , each loop the second summation need to start from zero so i guess to the second step is enough unless ur trying to prove something else , whats ur whole question anyway ?

OpenStudy (anonymous):

@wio @ikram002p I let \(r=j+k\) since the formula to prove was to get: \[\sum_{r=0}^{n+m}\left(\sum_{k=0}^ra_kb_{r-k}\right)x^r\] . I assumed this is how they got their \(r\) variable? So my problem was trying to go from where I left off in my last post. I tried drawing the points covered by the double summation here using \(m=5,n=2\) as an example (attached) Where the left drawing was for \( =\sum\limits_{k=0}^n\sum\limits_{r=k}^ma_kb_{r-k}x^r\) And the right drawing was for \(\sum\limits_{r=0}^{n+m}\left(\sum\limits_{k=0}^ra_kb_{r-k}\right)x^r\) I guess my understanding of the double summation where an index depends on the other is what is making me not understand ?